OpenStudy (anonymous):

270 degrees convert into radians

OpenStudy (anonymous):

4.71238898

OpenStudy (anonymous):

(270)*(pi) / 180 = 4.71 radians

OpenStudy (kropot72):

The exact conversion is as follows: There are pi radians in 180 degrees. There are pi/2 radians in 90 degrees. \[\large 90^{0}+ 180^{0}=270^{0}=\pi + \frac{\pi}{2}=\frac{3 \pi}{2}\ radians\]